OpenStudy (anonymous):

2 part question- help please! a. What is the maximum speed (in m/s) a vehicle (mass = 555 kg) can have at the top of a bump on a roller coaster (radius = 16.3 m) in order that it remain on the track? b.A pail of water is rotated in a vertical circle of radius 1.26 meters. What must be the minimum speed (in m/s) of the pail at the top of the circle in order for no water to spill out? Thanks so much!!!

OpenStudy (anonymous):

b) The forces on the bucket at the top of the loop (if it's going *just fast enough to keep the water in the bucket are \[\sum F_y = mg=ma_{centripetal}\] \[g=a_{centripetal}\] \[g=\frac{v^2}{r}\] With that you can solve for the velocity

OpenStudy (anonymous):

For part a, you need to set up the forces at the top of the track as \[ \sum F = N-mg=-ma_{cent}\] and the normal force has to be greater than zero for the car to still be on the track \[0<N=-ma+mg\] \[0<m(-a+g)\] \[g>a\] \[\frac{v^2}{r}<g\] The velocity has to be less than some value, so you solve for the max. velocity. B) Looking at our original expression for velocity, \[g=\frac{v^2}{r}\] In order for the water not to fall out of the bucket, its centripetal acceleration has to be *greater than or equal to* the gravity \[g≤\frac{v^2}{r}\] The velocity has to be greater than some value, so you solve for the minimum velocity.

OpenStudy (anonymous):

THANKS SO MUCH! this helped me so much! you rock.

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

Does it all make sense?

OpenStudy (anonymous):

yes very much so. thanks again!